ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 68 Deeba, Elias, University of Houston-Downtown, USA Learning Linear Algebra Using an Interactive Text This presentation will be based on the Interactive Linear Algebra with MapleV text that will be published by Springer-Verlag in their series Books in the Mathematical Sciences. The CD-ROM that accompanies the text includes : Lessons. Labs, Applications, Theory and a Library of linear algebra procedures written in Maple V that automate all linear algebra concepts in an introductory course. All these components of the text are supported by a Graphical User Interface (GUI) that makes it easy for the user to navigate among the different components of the text. We will describe how to use this interactive text in teaching a course in linear algebra. We will share with the participants some of the successes we have since implementing this text in teaching linear algebra. We will also demonstrate how we can implement similar approach to other course such as Calculus and Differential Equations.